OpenStudy (anonymous):

Hi have made an attempt to show that f=mv^2/R may be written as F=mw^2R using the following formulas v=2piRn/delta t , w=2pin/delta t and w= Delta theta/ delta t.. however i was unsuccessful please help

OpenStudy (jamesj):

Look at these two equations: \[ v = \frac{2 \pi R n}{\Delta t} \] \[ \omega = \frac{2 \pi n}{\Delta t} \] Notice that *immediately* you can write \[ v = \omega R \]

OpenStudy (anonymous):

thank you so much !!!!!!!